The probability that a student likes Math is 0.7 and the probability that a student likes History is 0.2 . If only 10% of students like both Math and History, what is the probability that a student chosen at random dislikes at least one of them?
Answer #1

Solution:

P(Maths)=P(M)=0.7

P(History)=P(H)=0.2

P(Math and History)=10%=P(M and H)=0.10

P(likes atleast one of them )=P(M)+P(H)-P(M and H)

=0.7+0.2-0.10

=0.80

probability that a student chosen at random dislikes at least one of them?

P(dislikes atleat one)=1-P(likes atleat one)

=1-0.80

=0.20

ANSWER:

he probability that a student chosen at random dislikes at least one of them=0.20